OpenStudy (anonymous):

Find the volume of the region above the triangle with vertices (0, 2), (1, 1), and (2, 2), and under the surface z = xy. If I want to integrate with y first, what should be the limit?

OpenStudy (anonymous):

should I be splitting the triangle into two parts and int with y?

OpenStudy (anonymous):

I meant MUST I.

OpenStudy (tkhunny):

1) Come to terms with your equations: y = 2 y = -x + 2 y = x 2) Decide which way to go. Either y = 2 to y = -x+2 for x between 0 and 1 and then y = 2 to y = x for x between 1 and 2 Or x = y to x = -y + 2 for y between 1 and 2 Let's see if you follow that far.

OpenStudy (anonymous):

I see. I've get to the ans. thanks!